# The earliest phases of CNO enrichment in galaxiesthanks: Full Tables 1 and 2 are only available at the CDS via anonymous ftp to cdsarc.u-strasbg.fr(130.79.128.5) or via http://cdsarc.u-strasbg.fr/viz-bin/cat/J/A+A/xxx/Ayyy

### v1

@@ 5.1 The origin of the scatter
Having calibrated our stochastic GCE model to reproduce the dispersion of our data sample through an appropriate choice of the dilution factors, we now delve into the nature of the polluters responsible for the variation of chemical abundance ratios within our stellar sample at low metallicity. We focus on [C/Fe] and [N/Fe], as for [O/Fe] we could obtain only upper limits in over half of our stellar sample, especially at $\rm[Fe/H]<-2.5$.
@@
In Figs. 5 to 7, we present ‘chemical density maps’ for the present-day stellar population in the MW halo, as predicted by our model. In all figures, different panels refer to different explosion energies of Pop III SNe contributing to the initial phases of chemical enrichment. As it is evident from the figures, different Pop III SN types contribute differently to the observed scatter. Low-energy Pop III SNe (faint and CCSNe) are responsible for high values of $\rm[C/Fe]>1$ and $\rm[N/Fe]>0.5$, and struggle to accurately reproduce the regions with $\rm[C/Fe]<0.5$ and $\rm[N/Fe]<0$. As the explosion energy of Pop III SNe increases (high-energy SNe and hypernovae), the $\rm[C/Fe]<0.5$ data points are well reproduced. In contrast, despite accounting for the influence of energetic Pop III SNe, regions at $\rm[N/Fe]<-0.5$ are sparsely recovered. This discrepancy does not necessarily indicate an issue, but rather suggests that the predominant polluters responsible for enriching these stars could be Pop II stars with diverse rotation velocities. At this point, it is essential to clarify that our study does not aim to identify the potential descendants of the first stars. Although our results seem indeed to suggest that among carbon- and nitrogen-enhanced stars there may be possible candidates, it is beyond the scope of this work to engage in such speculation. The main goal of this paper is to shed light onto the early evolution of our own and other galaxies, accounting for the contribution of all possible stellar polluters acting at early times.
@@
In the light of the comparison of our results for the MW with what can be observed at high redshift, it is important to discuss the density maps for $\rm log(N/O)$ vs $\rm log(O/H)+12$. These are presented in Fig. 7, although the presence of many upper/lower limits in our stellar sample prevents us from drawing definitive conclusions. As it is evident from the figure, our predictions cover a broad range in metallicity, $9.5\rm 7<log(O/H)+12<$ 9.5. Notably, there is a clear trend in the abundance of $\rm log(N/O)$ with varying energy levels of Pop III SNe. With pre-enrichment from faint SNe, it is possible to achieve extremely high values of $\rm log(N/O)$ up to $\approx-0.5$. As the energy of the Pop III SNe increases, regions with $\rm log(N/O)>-1$ become less populated, while regions with $\rm-2.5<log(N/O)<-1$ and $9.5\rm 8<log(O/H)<$ 9.5 start to fill, reflecting the influence of core-collapse and high-energy SNe. Finally, with hypernovae, areas with higher metallicity, $\rm log(O/H)>8.5$, and low values of $\rm log(N/O)<-2.5$ are well reproduced. It is important to highlight the effects of the two extreme Pop III SN energetic regimes: only faint SNe manage to populate the regions with $\rm log(N/O)>-1$, while hypernovae primarily impact regions with $\rm log(N/O)<-2.5$.
@@ 6 Early chemical enrichment: the gaseous medium of high-redshift galaxies
Recent JWST spectroscopy of high-redshift galaxies, particularly the most luminous ones, such as GN-z11 (Oesch et al., 2016) and CEERS-1019 (Finkelstein et al., 2017), has revealed gas substantially enriched in nitrogen (see Bunker et al., 2023; Cameron et al., 2023; Senchyna et al., 2023; Maiolino et al., 2023; Marques-Chaves et al., 2024; Senchyna et al., 2024), sparking renewed interest in the evolution of the CNO elements and prompting new inquiries into their origins. In this section, we aim to provide testable predictions for a MW-like galaxy observed at different redshifts, along with a possible interpretation of GN-z11, within the framework of our GCE model, including pre-enrichment by Pop III SNe with yields dependent on the mass, internal mixing level, and explosion energy of the progenitor star (Heger & Woosley, 2010).
@@ 6.1 MW counterparts at high redshift
Figure 8 also shows the $\rm log(N/O)$ values measured in GN-z11 at $z=10.6$ (Bunker et al., 2023; Cameron et al., 2023; Senchyna et al., 2023, 2024; Maiolino et al., 2023), in CEERS-1019 at $z=8.6$ (Marques-Chaves et al., 2024) and in RXCJ2248-ID at $z\sim 6$ (Topping et al., 2024). Note that while our predictions for the proto-MW successfully cover the observed metallicity range of $97<\rm log(O/H)+12<$ 9, they struggle to reproduce the high observed values of $\rm log(N/O)>-0.5$. This discrepancy is not unexpected: these systems are significantly brighter and more massive than what is expected for the MW progenitor at the same redshifts. For instance, at $z\approx 11$, the surface stellar density predicted for a proto-MW is $\Sigma_{\star}\sim 430$ M pc-2, whereas for GN-z11, it is approximately four times higher, $\Sigma_{\star}\sim 1650$ M pc-2 (Tacchella et al., 2023). On the other hand, our results indicate that it is possible to achieve values as high as $\rm log(N/O)>-0.5$ in some regions of a forming MW analog within the metallicity range of $8.0\rm 7.5<log(O/H)+12<$ 8.0. As already discussed in Sect. 5.1, such high N/O values are due to the contribution of faint Pop III SNe (see Fig. 7). Accounting for the effects of stellar rotation in $Z=0$ massive star models would likely lead to even higher N/O ratios (Meynet et al., 2006; Limongi & Chieffi, 2018), but these are not implemented in the yields of Heger & Woosley (2010) for Pop III stars adopted in this work.
Figure 8: Theoretical log(N/O) vs log(O/H)+12 for a proto-MW galaxy seen at different redshifts. Density maps and contours refer to the predictions of the stochastic GCE model, while the solid lines refer to the predictions of the homogeneous GCE model. Abundances measured in GN-z11 ($z=11$, Bunker et al., 2023; Cameron et al., 2023; Senchyna et al., 2023, 2024), CEER-1019 ($z=8.6$, Marques-Chaves et al., 2024), and RXCJ2248 ($z=6$, Topping et al., 2024) are also shown for comparison (symbols and shaded areas).
@@
Figure 9: Predicted log(N/O) vs log(O/H)+12 for GN-z11. Different evolutionary tracks represent different Pop III star-forming clumps, which merge into the main-branch of the galaxy. The tracks are colour-coded in time (left) and stellar mass (right). The light-purple trapezoid, the dark-purple rectangle and the grey stripe represent the conservative and fiducial results for the abundances of GN-z11 by Cameron et al. (2023). The blue points are the photoionisation modelling results for GN-z11 by Senchyna et al. (2023).
6.2 Interpreting GN-z11
@@
Figure 9 presents the elemental abundance ratio $\rm log(N/O)$ in the ISM of GN-z11 predicted by a model tailored to this galaxy (see Sect. 3.3) as a function of $\rm log(O/H)+12$, used as a proxy for metallicity. The different traces illustrate the chemical evolution of different clumps, each harbouring one Pop III polluter at the beginning of the simulation, at fixed density so as to reproduce the observed properties of GN-z11. In the left panel, the traces are colour-coded based on the time elapsed from the beginning of the simulation, whereas in the right panel different colours indicate different masses of the Pop III stars that pre-enriched the ISM. Note that we specifically selected models where the ISM was pre-enriched by faint Pop III SNe, as Pop III SNe with higher explosion energies are not able to achieve log(N/O) values that align with the observational data (see Fig. 7). We observe three distinct chemical enrichment paths, each directly linked to the mass of the Pop III SN that pre-enriched the ISM.
In the first path, the trace starts at low values of $\rm log(N/O)\lesssim-3$ and $7.57<\rm log(O/H)+12<$ 7.5, gradually increasing in $\rm log(N/O)$ over time. The starting point of the track reflects the enrichment from low-mass Pop III stars ($m_{\rm Pop\>III}\leq 20$ M), as shown in the right panel of Fig. 9, while the subsequent rapid increase in log(N/O) first and, then, also in $\rm log(O/H)+12$ is due to the subsequent contribution of massive Pop II stars.
The second path involves Pop III SNe with masses in the range 20–50 M that pre-enrich the ISM to very low metallicity levels, $\rm\log(O/H)<6.0$, and extremely high $\rm log(N/O)>0$ values. Following this initial phase, the $\rm log(N/O)$ ratio rapidly declines due to significant oxygen production from subsequent Pop II SNe, to slightly increase later mostly because of primary nitrogen production from Pop II massive stars boosted by rotation.
@@
Finally, the third path begins with pre-enrichment from massive Pop III SNe ($m_{\rm Pop\>III}>50$ M), resulting in high $\rm log(N/O)>-0.5$ values. This path shows an initial decreasing trend in metallicity with an approximately constant $\rm log(N/O)$ value, followed by a subsequent increase. The initial decline in $\rm log(O/H)$ is attributed to the dilution due to the infall of pristine gas, followed by a chemical enrichment path similar to the one of the previous scenario, predominantly driven by the contribution of successive generations of Pop II stars.
@@ 7.1 Milky Way galaxy
7.2 N-enhanced high-redshift systems
Different scenarios have been proposed to explain the extreme N/O abundance ratios measured in high-redshift systems, such as GN-z11 and CEERS-1019 (Cameron et al., 2023; Marques-Chaves et al., 2024). Kobayashi & Ferrara (2024) proposed a dual-burst model emphasizing the crucial role of Wolf-Rayet (WR) stars as N producers during the secondary burst phase. Prantzos et al. (2017) and Gratton et al. (2019) linked chemical peculiarities in globular clusters to the extreme patterns observed in galaxies like GN-z11. Charbonnel et al. (2023) and Nagele & Umeda (2023) explored the impact of metal-enriched stars, especially those within the mass range $10^{3}$$10^{5}$ M, on early galaxy chemistry. Isobe et al. (2023) investigated chemical abundances in 70 galaxies, correlating them with WR stars, supermassive stars (SMS), and tidal disruption events.

### v2

@@ 5.1 The origin of the scatter
Having calibrated our stochastic GCE model to reproduce the dispersion of our data sample through an appropriate choice of the dilution factors, we now delve into the nature of the polluters responsible for the variation of chemical abundance ratios within our stellar sample at low metallicity.
@@
In Figs. 5 to 9, we present ‘chemical density maps’ for the present-day stellar population in the MW halo, as predicted by our model. In all figures, different panels refer to different explosion energies of Pop III SNe contributing to the initial phases of chemical enrichment. As it is evident from the figures, different Pop III SN types contribute differently to the observed scatter. Low-energy Pop III SNe (faint and CCSNe) are responsible for high values of [C/Fe] $>1$, [N/Fe] $>0.5$, and [O/Fe] $>1.5$ and struggle to accurately reproduce the regions with [C/Fe] $<0.5$, [N/Fe] $<0$, and [O/Fe] $<1$. As the explosion energy of Pop III SNe increases (high-energy SNe and hypernovae), the [C/Fe] $<0.5$ and [O/Fe] $<1$ data points are well reproduced. In contrast, despite accounting for the influence of energetic Pop III SNe, regions at $\rm[N/Fe]<-0.5$ are sparsely recovered. This discrepancy does not necessarily indicate an issue, but rather suggests that the predominant polluters responsible for enriching these stars could be Pop II stars with diverse rotation velocities. At this point, it is essential to clarify that our study does not aim to identify the potential descendants of the first stars. Although our results seem indeed to suggest that among carbon-, nitrogen-, and oxygen-enhanced stars there may be possible candidates, it is beyond the scope of this work to engage in such speculation. The main goal of this paper is to shed light onto the early evolution of our own and other galaxies, accounting for the contribution of all possible stellar polluters acting at early times.
@@
In the light of the comparison of our results for the MW with what can be observed at high redshift, it is important to discuss the density maps for $\rm log(N/O)$ and $\rm log(C/O)$ vs $\rm log(O/H)+12$. These are presented in Figs. 8 and 9, although the presence of many upper/lower limits in our stellar sample prevents us from drawing definitive conclusions. As it is evident from the figures, our predictions cover a broad range in metallicity, $9.5\rm 7<log(O/H)+12<$ 9.5. Notably, there is a clear trend in the abundance of $\rm log(N/O)$ and $\rm log(C/O)$ with varying energy levels of Pop III SNe. With pre-enrichment from faint SNe, it is possible to achieve extremely high values of $\rm log(N/O)$ up to $\approx-0.5$ and $\rm log(C/O)$ up to 0.5 . As the energy of Pop III SNe increases, regions with $\rm log(N/O)>-1$ and $\rm log(C/O)>0$ become less populated, while regions with $\rm-2.5<log(N/O)<-1$ and $\rm-1<log(C/O)<0$ at $9.5\rm 7.5<log(O/H)<$ 9.5 start to fill, reflecting the influence of core-collapse and high-energy SNe. Finally, with hypernovae, areas with higher metallicity, $\rm log(O/H)>8.5$, and low values of $\rm log(N/O)<-2.5$ and $\rm log(C/O)<-1$ are well reproduced. It is important to highlight the effects of the two extreme Pop III SN energetic regimes: only faint SNe manage to populate the regions with $\rm log(N/O)>-1$ and $\rm log(C/O)>0$, while hypernovae primarily impact regions with $\rm log(N/O)<-2.5$ and $\rm log(C/O)<-1$
@@ 6 Early chemical enrichment: the gaseous medium of high-redshift galaxies
Figure 10: Theoretical log(N/O) vs log(O/H)+12 for a proto-MW galaxy seen at different redshifts. Density maps and contours refer to the predictions of the stochastic GCE model, while the solid lines refer to the predictions of the homogeneous GCE model. Abundances measured in GN-z11 ($z=11$, Bunker et al., 2023; Cameron et al., 2023; Senchyna et al., 2024), CEERS-1019 ($z=8.6$, Marques-Chaves et al., 2024), and RXCJ2248-ID ($z=6$, Topping et al., 2024) are also shown for comparison (symbols and shaded areas).
Recent JWST spectroscopy of high-redshift, metal-poor galaxies, particularly the most luminous ones, such as GN-z11 (Oesch et al., 2016) and CEERS-1019 (Finkelstein et al., 2017), has revealed gas substantially enriched in nitrogen (see Bunker et al., 2023; Cameron et al., 2023; Maiolino et al., 2023; Marques-Chaves et al., 2024; Senchyna et al., 2024), sparking renewed interest in the evolution of the CNO elements and prompting new inquiries into their origins. While N-emitters tend to have normal C content for their metallicities (see Schaerer et al., 2024, their figure 2), extreme C enhancement is found in a few systems (D’Eugenio et al., 2024), which deserves special attention. In this section, we aim to provide testable predictions for a MW-like galaxy observed at different redshifts, along with a possible interpretation of the high N/O (C/O) ratio measured for GN-z11 (GS-z12), within the framework of our GCE model, including pre-enrichment by Pop III SNe with yields dependent on the mass, internal mixing level, and explosion energy of the progenitor star (Heger & Woosley, 2010).
@@ 6.1 MW counterparts at high redshift
Figure 10 also shows the $\rm log(N/O)$ values measured in GN-z11 at $z=10.6$ (Bunker et al., 2023; Cameron et al., 2023; Maiolino et al., 2023; Senchyna et al., 2024), in CEERS-1019 at $z=8.6$ (Marques-Chaves et al., 2024) and in RXCJ2248-ID at $z\sim 6$ (Topping et al., 2024). Note that while our predictions for the proto-MW successfully cover the observed metallicity range of $97<\rm log(O/H)+12<$ 9, they struggle to reproduce the high observed values of $\rm log(N/O)>-0.5$. This discrepancy is not unexpected: these systems are significantly brighter and more massive than what is expected for the MW progenitor at the same redshifts. For instance, at $z\approx 11$, the surface stellar density predicted for a proto-MW is $\Sigma_{\star}\sim 430$ M pc-2, whereas for GN-z11, it is approximately four times higher, $\Sigma_{\star}\sim 1650$ M pc-2 (Tacchella et al., 2023). On the other hand, our results indicate that it is possible to achieve values as high as $\rm log(N/O)>-0.5$ in some regions of a forming MW analog within the metallicity range of $8.0\rm 7.5<log(O/H)+12<$ 8.0. As already discussed in Sect. 5.1, such high N/O values are due to the contribution of faint Pop III SNe (see Fig. 8). Accounting for the effects of stellar rotation in $Z=0$ massive star models would likely lead to even higher N/O ratios (Meynet et al., 2006; Limongi & Chieffi, 2018), but these are not implemented in the yields of Heger & Woosley (2010) for Pop III stars adopted in this work.
6.2 Extreme abundance ratios at high redshift
In this section, we deal with compact metal-poor, high-redshift systems that display C/O or N/O abundance ratios at variance with what is expected for their metallicities.
6.2.1 GN-z11
@@
Figure 11: Left and middle panels: predicted log(N/O) vs log(O/H)+12 for GN-z11. Different evolutionary tracks represent different Pop III star-forming clumps, which merge into the main-branch of the galaxy. The tracks are colour-coded in time elapsed since the beginning of star formation (left) and stellar mass (middle). The light-purple trapezoid, the dark-purple rectangle and the grey stripe represent the conservative and fiducial results for the abundances of GN-z11 by Cameron et al. (2023). The blue points are the photoionisation modelling results for GN-z11 by Senchyna et al. (2024). Right panel: star formation history of GN-z11 according to Tacchella et al. (2023, red line and shaded area) and as adopted in the model (grey line).
@@
Figure 11 presents the elemental abundance ratio $\rm log(N/O)$ as a function of $\rm log(O/H)+12$ as a proxy for metallicity in the ISM of GN-z11, predicted by a model tailored to this galaxy (see Sect. 3.3). The different traces illustrate the chemical evolution of different clumps, each harbouring one Pop III polluter at the beginning of the simulation, at fixed density so as to reproduce the observed properties of GN-z11. In the left panel, the traces are colour-coded based on the time elapsed from the beginning of star formation in the model, whereas in the middle panel different colours indicate different masses of the Pop III stars that pre-enriched the ISM. Note that we specifically selected models where the ISM was pre-enriched by faint Pop III SNe, as Pop III SNe with higher explosion energies are not able to achieve log(N/O) values that align with the observational data (see Fig. 8). The stellar mass and star formation rate (Fig. 11, right-hand panel) predicted by the model at the time of the observations, 440 Myr after the Big Bang, are $8.8\times 10^{8}$ M and 25 M yr-1, respectively, which perfectly align with the data (by construction).
Two distinct sets of evolutionary tracks are visible, each corresponding to different star formation bursts. The dark purple tracks in Figure 11 (left panel) represent the chemical evolution during the initial low-level star formation activity, which lasts for approximately 40 Myr from the onset of galaxy formation. The lighter-coloured tracks correspond to the second star formation burst, which begins 40 Myr after the beginning of the simulation and evolves for approximately 60 Myr. Both bursts are fed by infall of gas of primordial chemical composition and initiated by pre-enrichment from Pop III stars, which establishes the starting points of the $\rm log(N/O)$ vs $\rm log(O/H)+12$ tracks.
@@
In the first star-formation episode, we observe three distinct chemical enrichment paths, each directly linked to the mass of the Pop III SN that pre-enriched the ISM. In the first path, the trace starts at low values of $\rm log(N/O)\lesssim-1$ and $\rm log(O/H)+12<6.0$, gradually increasing in $\rm log(N/O)$ over time. The starting point of the track reflects the enrichment from low-mass Pop III stars ($m_{\rm Pop\>III}<40$ M), as shown in the middle panel of Fig. 11, while the subsequent increase in log(N/O) and slight decrease in $\rm log(O/H)+12$ is due at first to the continuous infall of pristine gas and then to the overwhelming contribution of massive Pop II stars. The second path involves Pop III SNe with masses in the range 40–80 M that pre-enrich the ISM to $\rm\log(O/H)+12<7.0$, and $\rm log(N/O)>-1.5$ values. Following this initial phase, the $\rm log(N/O)$ ratio rapidly declines due to significant oxygen production from subsequent Pop II SNe, to slightly increase later mostly because of primary nitrogen production from Pop II massive stars boosted by rotation. Finally, the third path begins with pre-enrichment from massive Pop III SNe ($m_{\rm Pop\>III}>80$ M), resulting in high $\rm log(N/O)\sim-1$ values and $7.5\rm 7.0<\log(O/H)+12<$ 7.5. This path shows an initial decreasing trend in metallicity (with an approximately constant $\rm log(N/O)$ value), followed by a subsequent increase. The initial decline in $\rm log(O/H)$ is attributed to the dilution due to the infall of pristine gas, followed by a chemical enrichment path similar to the one of the previous scenario, predominantly driven by the contribution of successive generations of Pop II stars.
As the second burst begins, fueled by infall of gas of primordial chemical composition, a second episode of pre-enrichment by Pop III stars occurs which, together with the cumulative enrichment from the first burst, establishes the starting point of the new evolutive track. Following this initial phase, the chemical evolution is predominantly driven by Pop II star formation, dominating the later stages of evolution. This path shows an initial decreasing trend in metallicity with an approximately constant $\rm log(N/O)$ value. The initial decline in $\rm log(O/H)$ is attributed to the dilution due to the infall of pristine gas, followed by chemical enrichment predominantly driven by the contribution of successive generations of Pop II stars.
Figure 12: Predicted log(C/O) vs log(O/H) + 12 for GS-z12. The evolutionary tracks represent different PopIII star-forming clumps that merge into the main branch of the galaxy. The tracks are colour-coded to indicate the time elapsed since the beginning of star formation (left panel) and the PopIII stellar mass (right panel). Also shown is the measured log(C/O) abundance for GS-z12, as derived by D’Eugenio et al. (2024). The blue marker represents the estimate obtained through a Bayesian modeling of the spectrum, while the grey stripe marks the lower limit derived using the photoionization models of Gutkin et al. (2016).
@@
6.2.2 GS-z12
Figure 12 shows the elemental abundance ratio $\rm log(C/O)$ as a function of $\rm log(O/H)+12$ in the ISM of GS-z12, based on predictions from a model specifically designed for this galaxy. In the case of GS-z12, a single star formation burst is considered (see Sect. 3.3). The evolutionary tracks are colour-coded similarly to those in Fig. 11, with the colours indicating either the time elapsed since the onset of star formation (Fig. 11, left panel) or the stellar mass of the Pop III progenitors (Fig. 11, right panel). Similar to what is found for GN-z11, distinct evolutionary tracks are clearly visible, primarily shaped by the masses of the Pop III stars that pre-enrich the ISM.
Three sets of evolutionary tracks can be recognized. In the first set, the tracks start at low values of $\rm log(C/O)\approx-0.5$ and $\rm log(O/H)+12<6.5$, showing a steady decrease of $\rm log(C/O)$ as time progresses. This initial phase is driven by enrichment from low-mass Pop III stars ($m_{\rm Pop:III}<20$ M) that produce significant amounts of carbon with minimal oxygen contribution. As Pop II stars enter the scene, $\rm log(C/O)$ decreases while the oxygen abundance increases, reflecting the dominant contribution of Pop II SNe during later evolutionary phases.
The second set of tracks involves a pre-enrichment by Pop III stars with masses between 40 and 60 M. These stars enrich the ISM to high metallicities, $\rm log(O/H)+12\sim 8.5$, and $-0.2\lesssim\rm log(C/O)\lesssim 0.1$. The metallicity initially decreases, while $\rm log(C/O)$ remains constant, due to the contribution of the primordial infalling gas, which dilutes the metals without affecting their abundance ratio. Subsequently, $\rm log(C/O)$ decreases due to the contribution to the enrichment by massive Pop II SNe.
The third path corresponds to Pop III stars with masses between 60 and 80 M. These massive stars pre-enrich the ISM to low metallicities, $\rm log(O/H)+12\lesssim 7.5$, while $\rm log(C/O)$ reaches high values, exceeding 0.25 dex. After an initial phase in which the infall of pristine gas dilutes all the metals in the same way, the ISM becomes enriched in oxygen from Pop II SNe and the $\rm log(C/O)$ ratio begins to decrease, while $\rm log(O/H)+12$ steadily increases.
The fourth evolutionary path is initiated by the most massive Pop III stars ($m_{\rm Pop:III}>80$ M), leading to initial $\rm log(C/O)\lesssim-0.4$ and high metallicity, $\rm log(O/H)\sim 8.5$. As the track progresses, there is a decrease in $\rm log(O/H)+12$, again due to dilution from infalling pristine gas, followed by an enrichment phase dominated by Pop II stars, which gradually increases the oxygen abundance. The $\rm log(C/O)$ values remain relatively stable, reflecting the fact that the most massive Pop III stars exploding as faint SNe are characterizes by C/O ratios in their ejecta that resemble those of Pop II CCSNe (see Fig. 9).
[…]
In conclusion, the elevated $\rm log(C/O)$ ratio of GS-z12 can be attributed to early carbon production by Pop III stars of masses 50–70 M, followed by later oxygen enrichment from Pop II stars.
@@ 7.1 Milky Way galaxy
7.2 C- and N-enhanced high-redshift systems
Different scenarios have been proposed to explain the extreme N/O abundance ratios measured in high-redshift systems, such as GN-z11 and CEERS-1019 (Cameron et al., 2023; Marques-Chaves et al., 2024). Kobayashi & Ferrara (2024) proposed a dual-burst model emphasizing the crucial role of Wolf-Rayet (WR) stars as N producers during the secondary burst phase. Prantzos et al. (2017) and Gratton et al. (2019) linked chemical peculiarities in globular clusters to the extreme patterns observed in galaxies like GN-z11. On the same line, D’Antona et al. (2023) proposed that the high N content of GN-z11 could be due to the formation of second-generation stars from pristine gas and asymptotic giant branch stellar ejecta in a massive proto-globular cluster. Charbonnel et al. (2023) and Nagele & Umeda (2023) explored the impact of metal-enriched stars, especially those within the mass range $10^{3}$$10^{5}$ M, on early galaxy chemistry. Isobe et al. (2023) investigated chemical abundances in 70 galaxies, correlating them with WR stars, supermassive stars (SMS), and tidal disruption events.